### v1

@@ Abstract
Reservoir Computing is an Unconventional Computation model to perform computation on various different substrates, such as RNNs or physical materials. The method takes a “black-box” approach, training only the outputs of the system it is built on. As such, evaluating the computational capacity of these systems can be challenging. We review and critique the evaluation methods used in the field of Reservoir Computing. We introduce a categorisation of benchmark tasks. We review multiple examples of benchmarks from the literature as applied to reservoir computing, and note their strengths and shortcomings. We suggest ways in which benchmarks and their uses may be improved to the benefit of the reservoir computing community.
@@
reservoir computing, benchmark
@@ 1 Introduction
There are two main variants. The Echo State Network (ESN) was first introduced by Jaeger (Jaeger,, 2010; Jaeger et al., 2007b, ), as a way of training Recurrent Neural Networks efficiently. A similar model, the Liquid State Machine (LSM), based on a spiking network model was later introduced by Maass (Maass et al.,, 2002), and was intended to present a more biologically plausible way of training Neural Networks.
@@ 1.2 Why Are Benchmarks Used?
Given benchmarks are tasks used in a specific way, are they indeed a good way of evaluating Reservoir Computing in general, or just for those specific tasks? This is not a commentary on the quality or effectiveness of individual benchmarks (although such commentary is made below) but a look at the purpose of their use. Another way of evaluating Reservoir Computers is CHARC (Dale et al.,, 2019), and there are other methods for evaluating generic task-independent properties such as Kernel Rank, Generalisation Rank, and Memory Capacity (sec.8). So why use problem-based benchmarks, specific to a single task, instead of the task-independent property measures like CHARC?
@@
Placing our work in the context of the literature is useful for several reasons: we can compare actual results, often (although not always) through the use of a common and well-described experimental procedure. One such example is the procedure of the spoken digits task, as first described in (Verstraeten et al.,, 2005). It can also help us showcase particular properties of a Reservoir Computer, such as the long-short term memory benchmarks described in (Jaeger,, 2012).
The quantifiable aspect is also useful: by being able to express the performance of a Reservoir Computer on a given benchmark, one can say that it is better or worse than others, as opposed to simply “different”.
@@ 1.3 Types of Reservoir Computing works
We divide the Reservoir Computing literature into three categories, similar to those described by Tanaka et al., (2019). Although not every reference falls cleanly into one or other of these, the categorisation provides a useful tool to identify what arguments a work is making, and which benchmarks may be best suited to support those arguments. We propose the following categories:
@@
These works typically focus on particular problems, and whether Reservoir Computing can provide a solution. An example is investigating the usefulness of Reservoir Computing when controlling robots (Plöger et al.,, 2003).
Works in this category typically do not use benchmarks, but they may still have a comparative element to them, such as comparing the Reservoir Computer’s performance to the current state of the art; for example, such as discrete Hidden Markov Models in investigations of speech recognition (Verstraeten et al.,, 2005; Triefenbach et al.,, 2010).
Although this category does not use benchmarks in service of the main argument, it is helpful to distinguish it, to provide a contrast to the categories that do, described below. It can also help to identify tasks that Reservoir Computing may be well suited to, such as Spoken Digit Recognition (Verstraeten et al.,, 2005), or predicting Spatio-Temporally Chaotic Systems like the Kuramoto–Sivashinsky Equation (Kuramoto,, 1978), that may subsequently become good benchmarks.
@@
Works in this category typically build on the original ESN (Jaeger,, 2010) and LSM (Maass et al.,, 2002) Reservoir Computing Models, in order to change or improve them. They may use benchmarks to demonstrate that their model is better than previous models; for example, where Jaeger (Jaeger et al., 2007a, ) introduces the concept of Leaky Integrator based ESNs. They may also use them to show that their model performs similarly to other models, but has other advantages, such as being simpler to implement in hardware. Examples of these include the introduction of the delay-line architecture (Appeltant et al.,, 2011), and in the analysis of reservoir topologies (Rodan and Tiňo,, 2010).
@@
These works are typically investigations into whether a particular material substrate is suitable as an implementation of a hardware Reservoir Computer, in terms of both performance and practicality. One of the first examples in this category, instantiating an LSM in a bucket of water (Fernando and Sojakka,, 2003), was used to illustrate properties of Reservoir Computers. The works in this category are reviewed by (Tanaka et al.,, 2019); examples include a reservoir instantiated in a swarm model (Lymburn et al.,, 2021), in a robot modelled on an octopus arm (Nakajima et al.,, 2013), and in carbon nanotubes (Dale et al.,, 2016).
1.4 Structure of this paper
@@
task type dynamics benchmark
@@
channel equalisation; two-jointed arm;
Van del Pol controller; pole balancing
@@
Mackey–Glass equation;
Multiple Superimposed Oscillators (MSO);
lazy eight figure; Lorenz chaos;
Kuramoto–Sivashinski equation
prediction unknown
sunspot numbers; Santa Fe laser data;
McMaster IPIX radar data
@@
MNIST handwritten digits
spoken digits; Japanese vowels;
Santa Fe sleap apnea data
@@
linear memory capacity (MC); nonlinear memory capacity;
kernel rank (KR); generalisation rank (GR)
[…]
@@ 2.2 Echo State Networks
Echo State Networks (fig. 1) are modelled as random recurrent neural networks, which have the form of a discrete time dynamical system. The vanilla form of an ESN is:
@@
Equation 1 defines the dynamics of the reservoir, where $\mathbf{x}$ is the $N$-d reservoir state; $\mathbf{W}$ is a random $N\times N$ weight matrix; $\mathbf{u}$ is the $N_{u}$-d input; $\mathbf{W}_{u}$ is a random input $N\times N_{u}$ weight matrix; and $f$ is a non-linear function (typically the tanh or other sigmoid function). Further terms may be included to allow leakiness, feedback, and so on.
Equation 2 defines the $N_{v}$-d output $\mathbf{v}$ as an observation of the reservoir state through a trained output $N_{v}\times N$ weight matrix $\mathbf{W}_{v}$. [^1] Some authors use $\mathbf{u}(t+1)$ and/or $\mathbf{v}(t)$ in their defining equations. See (Stepney,, 2021) for a discussion.
@@
An ESN is typically trained to match a target output $\hat{\mathbf{v}}$, often using ridge regression; a guide to this as well as a discussion of how to set various RC parameters can be found in Lukoševičius, (2012).
@@
  • In materio reservoirs (Tanaka et al.,, 2019) are physical reservoir computers, where the recurrent neural network model is (approximately) represented using some physical non-linear dynamical system.
  • Liquid State Machines (Fernando and Sojakka,, 2003) are a Reservoir Computing paradigm based on Spiking Neural Networks.
  • Delay-line reservoir (Appeltant et al.,, 2011) uses a single dynamical node, using a mask on the input such that it is multiplexed in time, rather than in space.
@@ 2.3.1 Dimensionality
The general reservoir equation (eqn. 1) allows for a vector-valued input at each timestep. Most of the benchmarks described here, such as NARMA (section 4.1) and the sunspot prediction benchmark (section 5.6), have a single-valued, or scalar-valued, input at each timestep ($N_{u}=1$). Some more complex multi-valued, or vector-valued, ($N_{u}>1$) input benchmarks, such as the XOR task (Fernando and Sojakka,, 2003), the Van der Pol Oscillator task (Gan,, 2023), and the sleep apnea task (section 7.4) are also available.
@@
A data set generated from an equation can be stationary, and has the advantage that the data set can be extended, simply by running the generating equation for longer. Gilpin, (2021) provides a catalogue of 131 chaotic dynamics systems equations, suitable for generating stationary time series for benchmarks. The catalogue includes the Mackey–Glass (section 5.1) and Lorenz systems (section 5.4).
@@ 2.3.3 Washout data
Typically, an initial subset of the data for a washout period. This puts the reservoir into a state so that results are independent of its earlier history, such as running a previous experiment, or of being switched off. This forgetting is possible because of the reservoir’s fading memory property.
@@ 2.4.2 Categorical encoding
In “one-hot” encoding, the reservoir has $N$ output nodes, one for each categorical value. At any given timestep, the output of the reservoir is taken to be the value of the node with the highest response (the one ‘hot’ node). As such, this encoding is suited to many classification tasks.
@@ 3.1 Main classes of benchmark tasks
We identify five main classes of benchmark task. The first two relate to training a reservoir to imitate an open (non-autonomous) dynamical system, or to predict an autonomous one. The next two classes relate to reservoirs performing specific computations on their inputs, or classifying their inputs. The final class is a set of measures of specific properties of reservoirs.
@@ 3.1.1 Imitation tasks
Figure 2: Training and testing stages of an imitation benchmark. The reservoir RC is fed the same time series input $\mathbf{u}$ as the target dynamical system DS. In the training stage, the output weights $\mathbf{W}_{v}$ of the reservoir are trained such that the resulting reservoir output $\mathbf{v}$ resembles the target dynamical system output $\hat{\mathbf{v}}$, the NRMSE (see section 9.3.1) can be used as an evaluation of the training. In the testing stage, the trained reservoir output weights are used, and we evaluate the reservoir using the error between the observed reservoir output $\mathbf{v}$ and the target dynamical system output $\hat{\mathbf{v}}$.
In an imitation task (see Figure 2), the Reservoir Computer is trained to replicate the dynamics of an open dynamical system. As such, it is given the same time series input as is given to the target dynamical system. This input is typically, but not necessarily, uniform random noise, such as with NARMA (section 4.1). The reservoir is then trained to produce the same time series output that the target system produces.
The performance of a Reservoir Computer on imitation tasks can be tested in two different ways. The first of these is to measure the error between the reservoir output values and the desired output values. In an imitation task, the desired output values are the output of the dynamical system being imitated. If the dynamical system being imitated is part of a greater whole, such as the control system of a robot, then another adequate test may be, “can the trained reservoir adequately replace the target system in situ”?
@@ 3.1.2 Prediction tasks
In a prediction task, the reservoir imitates a closed dynamical system. The reservoir must here predict the output of the dynamical system based on the previous output (figure 3). This is useful for learning the behaviour of a dynamical system where the input is unknown, or perhaps does not even exist. However, that is not a requirement. Any dynamical system that can be used for an imitation task can also be used for a prediction task: the only difference is the information provided to the reservoir.
@@
During training, the reservoir is fed outputs from the target dynamical system as inputs, an the output weights trained to minimise the difference. There are two options during the testing stage. In the driven case, the reservoir is driven with the target outputs as in the training stage. This provides a one-step look-ahead prediction, continually corrected by the actual system outputs to correct for drift. In the free-running case (Jaeger and Haas,, 2004) the reservoir is fed back its own previous output as its next input. This provides a multi-step look-ahead prediction, potentially subject to accumulating errors.
@@ 3.1.3 Computation tasks
In a computation task, the focus is on the reservoir performing logical or arithmetical computations on its inputs. Examples of logical operations include XOR (section 6.1), parity (section 6.2), and logic gates (Shougat et al.,, 2021). Examples of arithmetical operations are covered by other classifications: non-linear memory capacity (section 8.1.2) measures how well the reservoir can compute various polynomial functions of its inputs, and imitating or predicting systems with known dynamics (section 3.2.1) requires the reservoir to compute those dynamics.
@@ 3.1.4 Classification tasks
Speech recognition is a traditional classification benchmark task (Doddington and Schalk,, 1981). As such, it has also found a home as a Reservoir Computing benchmark, although not as widespread as the dynamical systems tasks, which tend to translate better to the intrinsically temporal Reservoir Computing paradigm.
@@
A style of classification task better suited to the medium of RC is time series based classification tasks, where the goal is to identify what the state of the time series is currently. This type of task has frequently been used in biomedical applications of RC, such as classifying gestures based on EMG signals (Donati et al.,, 2018; Garg et al.,, 2021), identifying arrhythmic heartbeats (Cucchi et al.,, 2021), or detecting seizures based on EEG signals (Kudithipudi et al.,, 2015). No standard RC benchmark task for this style of classification appears to have yet emerged, however. We suggest that the Santa Fe sleep apnea data set would be a suitable candidate for an RC classification benchmark task (section 7.4).
@@ 3.1.5 Property measures
Instead of measuring performance on a particular task, some benchmarks instead measure high level computational properties of the reservoir. The object of these tasks is not to maximise a success rate or minimise an error rate. Instead, we can take these properties and see how the behaviour of the reservoir may make them better suited to different tasks.
@@ 3.2 Other axes of classification
Are the dynamics of the benchmark task known and understood by us, or are they unknown, with only the outputs and perhaps the inputs being observed? In the first case, we typically have a dynamical system model that we can use to generate the behaviour, such as NARMA (section 4.1) or Mackey–Glass (section 5.1). The second case typically uses observed natural phenomena, like laser data (section 5.7) or sunspot observations (section 5.6), where we do not have a full model to generate the underlying dynamics, and are restricted to the supplied dataset.

### v2

@@ Abstract
Reservoir Computing is an Unconventional Computation model to perform computation on various different substrates, such as recurrent neural networks or physical materials. The method takes a ‘black-box’ approach, training only the outputs of the system it is built on. As such, evaluating the computational capacity of these systems can be challenging. We review and critique the evaluation methods used in the field of reservoir computing. We introduce a categorisation of benchmark tasks. We review multiple examples of benchmarks from the literature as applied to reservoir computing, and note their strengths and shortcomings. We suggest ways in which benchmarks and their uses may be improved to the benefit of the reservoir computing community.
@@
reservoir computing, benchmarks, NARMA, CHARC, memory capacity
@@ 1 Introduction
There are two main variants. The Echo State Network (ESN) was first introduced by Jaeger [1, 2, 3], as a way of training Recurrent Neural Networks efficiently. A similar model, the Liquid State Machine (LSM), based on a spiking network model was later introduced by Maass [4], and was intended to present a more biologically plausible way of training Neural Networks.
@@ 1.2 Why Are Benchmarks Used?
Given benchmarks are tasks used in a specific way, are they indeed a good way of evaluating Reservoir Computing in general, or just for those specific tasks? This is not a commentary on the quality or effectiveness of individual benchmarks (although such commentary is made below) but a look at the purpose of their use. Another way of evaluating Reservoir Computers is CHARC [5], and there are other methods for evaluating generic task-independent properties such as Kernel Rank, Generalisation Rank, and Memory Capacity (sec.8). So why use problem-based benchmarks, specific to a single task, instead of the task-independent property measures like CHARC?
@@
Placing our work in the context of the literature is useful for several reasons: we can compare actual results, often (although not always) through the use of a common and well-described experimental procedure. One such example is the procedure of the spoken digits task, as first described in [6]. It can also help us showcase particular properties of a Reservoir Computer, as do the long-short term memory benchmarks [7].
The quantifiable aspect is also useful: by being able to express the performance of a Reservoir Computer on a given benchmark, one can say that it is better or worse than others, as opposed to simply ‘different’.
@@ 1.3 Types of Reservoir Computing works
We divide the Reservoir Computing literature into three categories, similar to those described by [8]. Although not every reference falls cleanly into one or other of these, the categorisation provides a useful tool to identify what arguments a work is making, and which benchmarks may be best suited to support those arguments. We propose the following categories:
@@
These works typically focus on particular problems, and whether Reservoir Computing can provide a solution. An example is investigating the usefulness of Reservoir Computing when controlling robots [9].
Works in this category typically do not use benchmarks, but they may still have a comparative element to them, such as comparing the Reservoir Computer’s performance to the current state of the art; for example, such as discrete Hidden Markov Models in investigations of speech recognition [6, 10].
Although this category does not use benchmarks in service of the main argument, it is helpful to distinguish it, to provide a contrast to the categories that do, described below. It can also help to identify tasks that Reservoir Computing may be well suited to, such as Spoken Digit Recognition [6], or predicting spatio-temporally chaotic systems like the Kuramoto–Sivashinsky Equation [11], that may subsequently become good benchmarks.
@@
Works in this category typically build on the original ESN [1] and LSM [4] Reservoir Computing Models, in order to change or improve them. They may use benchmarks to demonstrate that their model is better than previous models; for example, where Jaeger [12] introduces the concept of Leaky Integrator based ESNs. They may also use them to show that their model performs similarly to other models, but has other advantages, such as being simpler to implement in hardware. Examples of these include the introduction of the delay-line architecture [13], and in the analysis of reservoir topologies [14].
@@
These works are typically investigations into whether a particular material substrate is suitable as an implementation of a hardware Reservoir Computer, in terms of both performance and practicality [15]. One of the first examples in this category, instantiating an LSM in a bucket of water [16], was used to illustrate properties of Reservoir Computers. The works in this category are reviewed by [8]; examples include a reservoir instantiated in a swarm model [17], in a robot modelled on an octopus arm [18], and in carbon nanotubes [19].
1.4 Structure of this review
Other authors provide overviews of a variety of RC benchmarks and problems, and some categorisation of tasks. For example, [20] provides a compact categorisation of many benchmarks and applications. Here, we take many of the more well-know, and well-used, benchmarks, delve into their details, variations, and limitations, and critique their use and applicability.
@@ 1.4 Structure of this review
task type dynamics benchmark section
@@
channel equalisation 4.2
two-jointed arm 4.3
Van del Pol controller 4.4
pole balancing 4.5
@@
Mackey–Glass equation 5.1
Multiple Superimposed Oscillators (MSO) 5.2
lazy figure eight 5.3
Lorenz chaos 5.4
Kuramoto–Sivashinski equation 5.5
unknown
sunspot numbers 5.6
Santa Fe laser data 5.7
McMaster IPIX radar data 5.8
@@
MNIST handwritten digits 7.1
spoken digits 7.2
Japanese vowels 7.3
Santa Fe sleep apnea data 7.4
@@
linear memory capacity (MC) 8.1.1
nonlinear memory capacity 8.1.2
kernel rank (KR) 8.2.1
generalisation rank (GR) 8.2.2
@@ 2.2 Echo State Networks
Echo State Networks (fig. 1) are modelled as random recurrent neural networks, which have the form of a discrete time dynamical system. The minimal form of an ESN is:
@@
Equation 1 defines the dynamics of the reservoir, where $\mathbf{x}$ is the $N$-d reservoir state; $\mathbf{W}$ is a random $N\times N$ weight matrix; $\mathbf{u}$ is the $N_{u}$-d input; $\mathbf{W}_{u}$ is a random input $N\times N_{u}$ weight matrix; and $f$ is a non-linear function (typically $\tanh(.)$ or other sigmoid function). Further terms may be included to allow leakiness, feedback, and so on.
Equation 2 defines the $N_{v}$-d output $\mathbf{v}$ as an observation of the reservoir state through a trained output $N_{v}\times N$ weight matrix $\mathbf{W}_{v}$. [^1] Some authors use $\mathbf{u}(t+1)$ and/or $\mathbf{v}(t)$ in their defining equations. See [21] for a discussion.
@@
An ESN is typically trained to match a target output $\hat{\mathbf{v}}$, often using ridge regression; a guide to this as well as a discussion of how to set various RC parameters can be found in [22].
@@
  • In materio reservoirs [8, 15] are a form of physical RC where the recurrent neural network model is (approximately) represented using some physical non-linear dynamical system.
  • Liquid State Machines [16] are a Reservoir Computing paradigm based on Spiking Neural Networks.
  • Delay-line reservoir [13] uses a single dynamical node, using a mask on the input such that it is multiplexed in time, rather than in space. These are commonly instantiated as a form of physical RC, but may also be simulated digitally.
@@ 2.3.1 Dimensionality
The general reservoir equation (eqn. 1) allows for a vector-valued input at each timestep. Most of the benchmarks described here, such as NARMA (section 4.1) and the sunspot prediction benchmark (section 5.6), have a single-valued, or scalar-valued, input at each timestep ($N_{u}=1$). Some more complex multi-valued, or vector-valued, ($N_{u}>1$) input benchmarks, such as the XOR task [16], the Van der Pol Oscillator task [23], and the sleep apnea task (section 7.4) are also available.
@@
A data set generated from an equation can be stationary, and has the advantage that the data set can be extended, simply by running the generating equation for longer. Gilpin [24] provides a catalogue of 131 chaotic dynamics systems equations, suitable for generating stationary time series for benchmarks. The catalogue includes the Mackey–Glass (section 5.1) and Lorenz systems (section 5.4).
@@ 2.3.3 Washout data
Typically, an initial subset of the data is used for a washout period. This puts the reservoir into a state so that results are independent of its earlier history, such as running a previous experiment, or of being switched off. This forgetting is possible because of the reservoir’s fading memory property.
@@ 2.4.2 Categorical encoding
In ‘one-hot’ encoding, the reservoir has $N$ output nodes, one for each categorical value. At any given timestep, the output of the reservoir is taken to be the value of the node with the highest response (the one ‘hot’ node). As such, this encoding is suited to many classification tasks.
@@ 3.1 Main classes of benchmark tasks
We identify five main classes of benchmark task. The first two relate to training a reservoir to imitate an open (non-autonomous) dynamical system, or to predict a closed (autonomous) one. The next two classes relate to reservoirs performing specific computations on their inputs, or classifying their inputs. The final class is a set of measures of specific properties of reservoirs. As the subject develops, and RCs are applied to wider classes of problems, new classes of task may be devised.
@@ 3.1.1 Imitation tasks
Figure 2: Training and testing stages of an imitation benchmark. The reservoir RC is fed the same time series input $\mathbf{u}$ as the target dynamical system DS. In the training stage, the output weights $\mathbf{W}_{v}$ of the reservoir are trained such that the resulting reservoir output $\mathbf{v}$ resembles the target dynamical system output $\hat{\mathbf{v}}$, the Normalised Root Mean Square Error (NRMSE, see section 9.3.1) can be used as an evaluation of the training. In the testing stage, the trained reservoir output weights are used, and we evaluate the reservoir using the error between the observed reservoir output $\mathbf{v}$ and the target dynamical system output $\hat{\mathbf{v}}$.
In an imitation task (see Figure 2), the Reservoir Computer is trained to replicate the dynamics of an open dynamical system, one with inputs. As such, it is given the same time series input as is given to the target dynamical system. This input may, for example, be uniform random noise, such as with NARMA (section 4.1), or a regular frequency signal, such as used in the classic sine wave generation task. The reservoir is then trained to produce the same time series output that the target system produces.
The performance of a Reservoir Computer on imitation tasks can be tested in two different ways. The first of these is to measure the error between the reservoir output values and the desired output values. In an imitation task, the desired output values are the output of the dynamical system being imitated. If the dynamical system being imitated is part of a greater whole, such as the control system of a robot, then another adequate test may be, ‘can the trained reservoir adequately replace the target system in situ’?
@@ 3.1.2 Prediction tasks
In a prediction task, the reservoir imitates a closed dynamical system, one with no inputs. The reservoir must here predict the output of the dynamical system based on the previous output (figure 3). This is useful for learning the behaviour of a dynamical system where the input is unknown, or perhaps does not even exist. However, that is not a requirement. Any dynamical system that can be used for an imitation task can also be used for a prediction task: the only difference is the information input to the reservoir.
@@
During training, the reservoir is fed outputs from the target dynamical system as inputs, an the output weights trained to minimise the difference. There are two options during the testing stage. In the driven case, the reservoir is driven with the target outputs as in the training stage. This provides a one-step look-ahead prediction, continually corrected by the actual system outputs to correct for drift. In the free-running case [25] the reservoir is fed back its own previous output as its next input. This provides a multi-step look-ahead prediction, potentially subject to accumulating errors.
@@ 3.1.3 Computation tasks
In a computation task, the focus is on the reservoir performing specific logical or arithmetical computations on its inputs. Examples of logical operations include XOR (section 6.1), parity (section 6.2), and logic gates [26]. Examples of arithmetical operations are covered by other classifications: non-linear memory capacity (section 8.1.2) measures how well the reservoir can compute various polynomial functions of its inputs, and imitating or predicting systems with known dynamics (section 3.2.1) requires the reservoir to compute those dynamics.
@@ 3.1.4 Classification tasks
Speech recognition is a traditional classification benchmark task [27]. As such, it has also found a home as a Reservoir Computing benchmark, although not as widespread as the dynamical systems tasks, which tend to translate better to the intrinsically temporal Reservoir Computing paradigm.
@@
A style of classification task better suited to the medium of RC is time series based classification tasks, where the goal is to identify what the state of the time series is currently. This type of task has frequently been used in biomedical applications of RC, such as classifying gestures based on EMG signals [28, 29], identifying arrhythmic heartbeats [30], or detecting seizures based on EEG signals [31]. No standard RC benchmark task for this style of classification appears to have yet emerged, however. We suggest that the Santa Fe sleep apnea data set would be a suitable candidate for an RC classification benchmark task (section 7.4).
@@ 3.1.5 Property measures
Rather than measuring performance on a particular task, some benchmarks instead measure more direct computational properties of the reservoir, such as its memory and generalisation capabilities. The object of these tasks is not to maximise a success rate or minimise an error rate. Instead, we can take these properties and see how the behaviour of the reservoir may make them better suited to different tasks.
@@ 3.2 Other axes of classification
Are the dynamics of the benchmark task known and understood by us, or are they unknown, with only the outputs and perhaps the inputs being observed? In the first case, we typically have a dynamical system model that we can use to generate the behaviour, such as NARMA (section 4.1) or Mackey–Glass (section 5.1). The second case typically uses observed natural phenomena, like laser data (section 5.7) or sunspot observations (section 5.6), where we do not have a full model to generate the underlying dynamics, and are restricted to the supplied real world dataset.